OpenStudy (anonymous):

Solve the following system of equations. 2x - 3y + z = -1 2x - 2y - 2z = -8 x + 5y - 3z = 2

OpenStudy (texaschic101):

2x - 3y + z = -1 2x - 2y - 2z = -8 -->(-1) ----------------- 2x - 3y + z = -1 -2x + 2y + 2z = 8 (result of multiplying by -1) ----------------add -y + 3z = 7 2x - 2y - 2z = -8 x + 5y - 3z = 2 -->(-2) --------------- 2x - 2y - 2z = -8 -2x - 10y + 6z = -4 (result of multiplying by -2) ---------------add -12y + 4z = - 12 -y + 3z = 7 -->(-12) -12y + 4z = -12 ---------------- 12y - 36z = - 84 (result of multiplying by -12) -12y + 4z = -12 ----------------add - 32z = - 96 z = -96/-32 z = 3 -y + 3z = 7 -y + 3(3) = 7 -y + 9 = 7 -y = 7 - 9 -y = - 2 y = 2 2x - 3y + z = -1 2x - 3(2) + 3 = -1 2x - 6 + 3 = -1 2x - 3 = -1 2x = -1 + 3 2x = 2 x = 1 check... x + 5y - 3z = 2 1 + 5(2) - 3(3) = 2 1 + 10 - 9 = 2 11 - 9 = 2 2 = 2 (correct) x = 1, y = 2, and z = 3

OpenStudy (texaschic101):

its not really that hard...just time consuming. Do you have any questions ?